OpenStudy (anonymous):

If (x+1)(y-3), and x is positive, what is the value of y

OpenStudy (anonymous):

incomplete question

OpenStudy (anonymous):

well thats what is says on my homework. xP

OpenStudy (anonymous):

i think it has to do with factoring

OpenStudy (mertsj):

You need to post the ENTIRE question EXACTLY as written.

OpenStudy (anonymous):

thats ALL thats written. except it gives some answer choices.

OpenStudy (mertsj):

You need to post the ENTIRE question EXACTLY as written.

OpenStudy (anonymous):

BUT THATS WHATS WRITTEN. OK. theres NOTHING ELSE :(

OpenStudy (mertsj):

except it gives some answer choices.

OpenStudy (anonymous):

yes. and would you like to see the answer choices.

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

(x+1)(y-3=0 A - 2 B- 6 C-3 D- 6 E- 22

OpenStudy (anonymous):

ok if \(x\) is positive, then \(x+1\) is also positive and therefore not equal to zero therefore, if \((x+1)(y-3)=0\) since \(x+1\neq 0\) you must have \(y-3=0\) making \(y=3\)

OpenStudy (anonymous):

the incomplete part was the missing \(=0\) in the question

OpenStudy (anonymous):

ohh! thank you so much!